OpenStudy (clamin):

PLEASE HELP!! MEDAL!! Simplify the expression: k+1/k^2-6k-7

Miracrown (miracrown):

\[\frac{ K + 1 }{ k ^{2} - 6k - 7}\]

Miracrown (miracrown):

We can first factor the denominator

OpenStudy (clamin):

k+1/(k-7)(k+1)

OpenStudy (clamin):

@Miracrown

Miracrown (miracrown):

Yes, that's right

Miracrown (miracrown):

And we can cancel out k+1

Miracrown (miracrown):

So, after cancelling k+1, what does it become?

OpenStudy (clamin):

k-7??

Miracrown (miracrown):

Nope, it becomes 1/k-7

OpenStudy (clamin):

why it becomes 1??

Miracrown (miracrown):

k+1=1*(k+1) k+1 is canceled out, but 1 is still there

OpenStudy (clamin):

i have another question is 3-x = 3??

Miracrown (miracrown):

No. What made you think that?

Miracrown (miracrown):

subtract 3 on both sides -x=0 so x=0

OpenStudy (clamin):

idk..its just i did the same thing as you did for cancelling out the k+1 and becomes 1

OpenStudy (clamin):

x^2-4x+3/3-x

OpenStudy (clamin):

@Miracrown

Miracrown (miracrown):

3-x=-(x-3)

Miracrown (miracrown):

Once you cancel out x-3, the denominator becomes -1